1. Students are asked to develop a plan for separating Al2O3(s) from a mixture of powdered Fe2O3(s) and powdered Al2O3(s) using chemical reactions and laboratory techniques.
  2. One student proposes that Al2O3(s) can be separated from the mixture by adding water to the mixture and then filtering. Explain why this approach is not reasonable.
  3. A second student organizes a plan using a table. The first two steps have already been entered in the table as shown in the upper table above. Complete the plan by listing the additional steps that are needed to recover the Al2O3(s). List the steps in the correct order and refer to the appropriate reaction by number, if applicable.
  4. The second student recovers 5.5 g of Al2O3(s) from a 10.0 g sample of the mixture. Calculate the percent of Al by mass in the mixture of the two powdered oxides. (The molar mass of Al2O3 is 101.96 g/mol, and the molar mass of Fe2O3 is 159.70 g/mol.)
1
Expert's answer
2021-04-30T04:52:25-0400

2) These oxides do not interact with water (do not dissolve in it).


3) Table data is required.


4)

"n(Al_2O_3) = m\/M = 5.5\/101.96 = 0.054 mol"

"n(Al) = 2*0.054 = 0.108 mol"

"m(Al) =0.108*26.98= 2.9g"

"w(Al) = 2.9\/10.0 = 0.29"